Coastal Shores Inc. (CSI) was destroyed by Hurricane Fred on August 5, 2021. At January 1, CSI reported an inventory of $171,000. Sales from January 1, 2021, to August 5, 2021, totaled $481,000 and purchases totaled $196,000 during that time. CSI consistently marks up its products 65% over cost to arrive at a selling price. The estimated inventory loss due to Hurricane Fred would be:

  • $141,575.

  • $75,485.

  • $79,485.

  • None of these answer choices are correct.
Answer #1

Cost of Goods available for sale = Inventory + Purchases

= 171,000 + 196,000 = 367,000

Cost of Goods Sold = 481,000/165% = 291,515

Goods lost = Cost of Goods available for sale - Cost of Goods Sold

= 367,000 - 291,515

= 75,485